OpenStudy (rock_mit182):

Suppose that the functions and are defined as follows. find (q ° r) and (r ° q)

OpenStudy (rock_mit182):

\[q(x) = x ^{2} + 3\]\[r(x) = \sqrt{x+2}\]

OpenStudy (rock_mit182):

(q ° r )(2) = 3 (r ° q)(2) = 7

OpenStudy (rock_mit182):

is it right ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[q\circ r(2)=q(r(2))\] first \(r(2)=\sqrt{2+2}=\sqrt4=2\) and then \[q(r(2))=q(2)=2^2+3\] i think you had them backwards

OpenStudy (rock_mit182):

yeah i had, fortunally i realized before answer it